OpenStudy (anonymous):

| 4 - 2x | = 14 can you explain how to do this?

OpenStudy (anonymous):

yes. you want to get rid of the absolute values signs so you can solve a linear equation. the way you do that is to write \[4-2x=14\] OR \[4-2x=-14\] since the absolute value of both 14 and -14 is 14

OpenStudy (anonymous):

then you solve the two linear equations separately and get two possible answers

OpenStudy (anonymous):

if you are like me and would prefer to see \[2x-4=14\] and \[2x-4=-14\] you can do that as well. with a little experience you will see that \[|a-b|=|b-a|\] so you are permitted to change \[|4-2x|=|2x-4|\]of course it makes no difference either way will work

OpenStudy (anonymous):

ok I think I get it

OpenStudy (anonymous):

x = 3.5, -7?

OpenStudy (anonymous):

do just what they told you to do. and uyoull get it right every time.